# Emergent Fracton Hydrodynamics in the Fractional Quantum Hall Regime of Ultracold Atoms

### v1

@@ IV Fractional Quantum Hall Physics with Ultracold Atoms
We will now discuss the experimental requirements for probing fracton hydrodynamics with ultracold atoms subjected to synthetic gauge fields, focusing on the one hand on rotating quantum gases in the continuum in Sec. IV.1 and on the other hand ultracold atoms in optical lattices in Sec. IV.2. The squeezing terms of the effective one-dimensional models projected to the LLL are expressed in momentum basis, see Eqs. (1) and (5). However, both in the continuum and on the lattice, wave functions corresponding to different momenta $q$ in the LLL are the same apart for a translation in the $x$-direction (which we disucuss for the lattice in App. D, while for the continuum this can be directly read off from the wave function). As a consequence, we can equally interpret the projected Hamiltonians to describe squeezing motion in $x$-direction of real space and, hence, the dipole moment is also conserved in real space along the $x$-direction $\mathcal{P}=\sum_{x,y}xn_{x,y}$. We will now further elaborate on this point and analyze the full unprojected dynamics. This will allow us further to determine the minimal requirements for observing squeezing dynamics for both rotating quantum gases and ultracold atoms in optical lattices.
@@ IV.1 Rotating quantum gases in the continuum
In order to gain some insight into the dynamics of particles prepared in the LLL, we can consider a minimal model of two particles in a magnetic field, $H=H_{1}+H_{2}=\frac{1}{2}\big{[}(\mathbb{p}_{1}-\mathbb{a}_{1})^{2}+(\mathbb{p}_{2}-\mathbb{a}_{2})^{2}\big{]}$ where $\mathbb{a}_{1,2}=(0,x_{1,2})$ in the Landau gauge. The Hamiltonian is separable in center of mass $\mathbb{P},\mathbb{R}$ and relative $\mathbb{p},\mathbb{r}$ degrees of freedom, $H=\frac{1}{2}\big{[}\frac{1}{2}(\mathbb{P}-2\tilde{\mathbb{A}})^{2}+2(\mathbb{p}-\frac{1}{2}\mathbb{a})^{2}\big{]}$, where $\tilde{\mathbb{A}}=(0,X)$ and $\mathbb{a}=(0,x)$, with spectrum $E_{n,m}=\hbar\omega_{c}(n+m+1)$.
It follows that, when the initial state is fully contained in the LLL, both the center-of-mass wave function and relative distance wave function will be eigenstates in the LLL. For density-density interactions only the relative distance wave function evolves, while the center-of-mass wave function is still an eigenstate of the Hamiltonian. Thus, both average center of mass and its higher-moments are constant, and the dipole moment in real space is preserved to all orders. This reasoning can be directly generalized to more particles as well.
@@
When the center-of-mass wave function of the initial state is not an eigenstate, i.e., when the initial state of the center of mass is not confined in the LLL, higher moments of the dipole moment will not be conserved in general. We consider an initial state that is a superposition of different Landau levels both in the center of mass and relative coordinate $\Psi(\mathbb{R},\mathbb{r})=\sum_{\gamma\beta,n,m}c^{\gamma,\beta}_{n,m}\psi^{\gamma}_{n}(\mathbb{R})\phi^{\beta}_{m}(\mathbb{r})$, where $\psi(\mathbb{R})$ and $\phi(\mathbb{r})$ are eigenstates of the non-interacting problem for the center of mass and relative coordinate, respectively. Specifically they are given by
$\psi_{n}^{\gamma}(\mathbb{R})=\frac{1}{\sqrt{{\pi}n!2^{n}L_{y}}}H_{n}\bigg{(}\frac{X-X_{\gamma}}{L_{B}}\bigg{)}e^{-\frac{1}{2}(\frac{X-X_{\gamma}}{L_{b}})^{2}}e^{iK_{\gamma}Y}$ (12)
@@
where $K_{\gamma}=2\pi\gamma/L_{y}$, $X_{\gamma}=L_{b}^{2}K_{\gamma}$, and the magnetic length for the center of mass coordinate $L_{b}=\frac{l_{b}}{\sqrt{2}}$, as well as,
$\phi_{m}^{\beta}(\mathbb{r})=\frac{1}{\sqrt{{\pi}m!2^{m}L_{y}}}H_{m}\bigg{(}\frac{x-x_{\beta}}{\ell_{b}}\bigg{)}e^{-\frac{1}{2}(\frac{x-x_{\beta}}{\ell_{b}})^{2}}e^{ik_{\beta}y}$ (13)
@@
where $k_{\beta}=2\pi\beta/L_{y}$, $x_{\beta}=\ell_{b}^{2}k_{\beta}$, and the magnetic length for the relative coordinate $\ell_{b}=l_{b}\sqrt{2}$. By using the properties of the Hermite polynomials $H_{n}$ we determine the evolution of the moments of the center-of-mass in real-space in the $x$-direction (i.e., dipole moment in the $x$-direction in real space) $\mathcal{P}=X$ and find that for general initial states the center of mass is not conserved.
@@
As a concrete example, we consider an initial state where the two particles have momenta $k_{y}=0$ and that is invariant under inversion symmetry in the spatial coordinate. Since the two particles are bosons the wave function $\Psi(\mathbb{R},\mathbb{r})=\sum_{\alpha,\beta,n,m}c^{\gamma,\beta}_{n,m}\psi^{\gamma}_{n}(\mathbb{R})\phi^{\beta}_{m}(\mathbb{r})$, where $\psi(\mathbb{R})$ and $\phi(\mathbb{r})$ must be invariant under exchange of their coordinates $\Psi(\mathbb{R},-\mathbb{r})=\Psi(\mathbb{R},\mathbb{r})$, which implies $m$ must be even. Considering the state to be invariant under inversion symmetry, $\Psi(-\mathbb{R},-\mathbb{r})=\Psi(\mathbb{R},\mathbb{r})$, $n$ is even as well. Under these assumptions, the average position of the center of mass does not evolve. For concreteness, we show the evolution of the initial state $\Psi_{0}(\mathbb{R},\mathbb{r})=c^{0,0}_{0,0}\psi^{0}_{0}(\mathbb{R})\phi^{0}_{0}(\mathbb{r})+c^{0,0}_{2,2}\psi^{0}_{2}(\mathbb{R})\phi^{0}_{2}(\mathbb{r})$, for different ratios of $|c^{0,0}_{2,2}/c^{0,0}_{0,0}|$ in Fig 4. The fluctuations of the center of mass oscillate with frequency $2\omega_{c}$ when multiple Landau levels are occupied. When the initial state is fully contained in the LLL, $c^{0,0}_{2,2}=0$, the fluctuations of the dipole-moment $\langle\Delta\mathcal{P}^{2}\rangle$ (as well as all higher moments) are constant as argued before.
@@ IV.2 Synthetic gauge fields for ultracold atoms in optical lattices
Artificial gauge fields for charge neutral atoms in optical lattices can be created by different approaches. One route is based on using a set of nuclear spin states to realize a synthetic dimension within an atom. Tunneling matrix elements between these nuclear spin states with level dependent phase factors are realized by a two-photon Raman coupling, which realizes an artificial magnetic flux [19, 20, 56]. Another route is to use Floquet engineering to imprint the relevant phases on the tunneling matrix elements [13, 14, 15, 16, 17, 18, 21]. The resulting interacting lattice systems with flux are effectively described by the Hubbard-Hofstadter model, Eq. (2), and geometric conditions on the lattice can be defined such as the non-interacting bands are flat.
@@
We now investigate the dynamics of the full two-dimensional Hubbard-Hofstadter model of Eq. (2). To this end, we initialize the system at time $t=0$ with particles occupying orbitals of the lowest band of the non-interacting problem $H_{0}$. Experimentally, such a state can be obtained by preparing plane-wave states in the $y$-direction, corresponding to an eigenstate of the non-interacting Hofstadter model with ${t_{x}}/{t_{y}}\to 0$, followed by a slow increase of ${t_{x}}/{t_{y}}$; c.f. Ref. [23] where a similar approach has been considered. At time $t=0$ the on-site interaction energy is turned on and the particles start to evolve.
@@
Contrary to the continuum case, the Hamiltonian is not separable in center of mass and relative coordinate degrees of freedom. Thus, although the states exhibit an average conservation of the center of mass along the $x$-direction, $\mathcal{P}=\sum_{x,y}x\,n_{x,y}$, when their density profile is invariant under inversion symmetry (not shown), the fluctuations $\langle\Delta^{2}\mathcal{P}\rangle=\langle\mathcal{P}^{2}\rangle-\langle\mathcal{P}\rangle^{2}$ will not be constant in the presence of interband transitions (even when initially prepared in the LLL). On the other hand, when the dynamics is projected in the LLL, the value of the fluctuations (as well as higher momenta of the center of mass) is preserved. This is because similarly to the continuum the shape of the wave functions in the LLL remain invariant for different momenta and only their wave function center shifts along the $x$-direction, see App. D. Hence, a squeezing motion of the column density is expected in real space provided the dynamics is confined to the LLL. This formally follows also from the fact that the center of mass along the $x$-direction $\mathcal{P}$ projected onto the LLL, commutes with the projected interactions, see App. D. By contrast, for interlevel processes $\mathcal{P}$ is not conserved in real space, despite the total momentum $\sum k_{y}n_{k_{y}}$ being conserved, where the occupation of momentum $k_{y}$, $n_{k_{y}}=\sum_{m}n_{k_{y},m}$, is obtained by summing over all Landau levels $m$. Eigenfunctions with the same $k_{y}$ but on different Landau levels thus have different expectation values of the dipole-moment $\mathcal{P}$ evaluated in real space. We compute numerically the fluctuations of the dipole moment $\langle\Delta^{2}\mathcal{P}\rangle=\langle\mathcal{P}^{2}\rangle-\langle\mathcal{P}\rangle^{2}$ and find that they are reduced with decreasing interactions as the dynamics becomes increasingly confined to the LLL, see Fig. 5(d), consistent with the discussion above.
@@ V Discussion and Outlook
Recent experiments with ultracold atoms have realized models with synthetic gauge fields. In these systems, non-equilibrium quantum dynamics are most naturally accessible. Here, we have shown that the far-from-equilbrium dynamics in the LLL are governed by fracton hydrodynamics, characterized by a subdiffusively slow relaxation of density excitations and fluctuations. The properties of the relaxation are tunable by the effective thickness of the torus. The equivalence of dipole conservation in the LLL and dipole conservation in real space along the $x$-direction allows for a direct measurement of the unconventional relaxation dynamics, both in rotating quantum gases and in Hubbard-Hofstadter models with certain geometric constraints.
For future work, it will be interesting to analyze the short time dynamics of specific initial states in more detail. Furthermore, one could study the consequences of non-flat bands in the Hubbard-Hofstadter model, that arise for example with open boundary conditions, on the effective relaxation dynamics. Moreover, excitations of the FQH system could be studied in the low-temperature regime and related to the different dynamical regimes of fracton excitations [57, 58]. It would be also exciting to explore relaxation processes in the vortex dynamics of rotating quantum gases [6]. Previous works have already interpreted vortices as fractons and discussed Hilbert-space fragmentation phenomena and fractonic dynamics in vortex systems [59].
@@ Appendix B Generalization to different geometries and non-flat bands
The analysis in the main text focuses on the flat band condition, obtained directly for rotating quantum gases or for specific choices of lattice geometries in the Harper-Hofstadter model. Here, we generalize the considerations for lattice systems to non-flat bands. Non flat-bands arise in a lattice when the condition $N_{x}=Q=\alpha N_{y}$, is not satisfied and when the boundary conditions are not met.

### v2

@@ IV Fractional Quantum Hall Physics with Ultracold Atoms
Here, we explore how the dynamics is affected by considering all Landau levels and investigate the role of inter-level transitions. We particularly focus on two types of experimental platforms: rotating quantum gases (Sec. IV.1) and ultracold atoms in optical lattices (Sec. IV.2); and consider the full unprojected continuum and lattice models. Our analysis also clarifies the connection between the effective projected one-dimensional lattice and real space.
The squeezing terms of the effective one-dimensional models projected to the LLL are expressed in momentum basis, see Eqs. (1) and (5). However, both in the continuum and on the lattice, wave functions corresponding to different momenta $q$ in the LLL are the same apart for a translation in the $x$-direction (which we discuss for the lattice in App. D, while for the continuum this can be directly read off from the wave function). As a consequence, we can equally interpret the projected Hamiltonians to describe squeezing motion in $x$-direction of real space and, hence, the dipole moment is also conserved in real space along the $x$-direction $\mathcal{P}=\sum_{x,y}xn_{x,y}$. The aim of this section is to analyze the minimal requirements for observing squeezing dynamics in real space for both rotating quantum gases and ultracold atoms in optical lattices, when the full unprojected dynamics is considered. The initial states of the dynamics here considered are states where all the particles are initialized in a finite energy density state in the LLL, which will give rise to a similar relaxation dynamics to the ones considered in the previous section.
@@ IV.1 Rotating quantum gases in the continuum
In order to gain some insight into the dynamics of particles prepared in the LLL, we can consider a minimal model of two particles in a magnetic field, $H=H_{1}+H_{2}=\frac{1}{2}\big{[}(\mathbb{p}_{1}-\mathbb{a}_{1})^{2}+(\mathbb{p}_{2}-\mathbb{a}_{2})^{2}\big{]}$ where $\mathbb{a}_{1,2}=(0,x_{1,2})$ in the Landau gauge. The Hamiltonian is separable in center of mass $\mathbb{P},\mathbb{R}$ and relative $\mathbb{p},\mathbb{r}$ degrees of freedom, $H=\frac{1}{2}\big{[}\frac{1}{2}(\mathbb{P}-2\tilde{\mathbb{A}})^{2}+2(\mathbb{p}-\frac{1}{2}\mathbb{a})^{2}\big{]}$, where $\tilde{\mathbb{A}}=(0,X)$ and $\mathbb{a}=(0,x)$, with spectrum $E_{\gamma_{1},\gamma_{2}}=\hbar\omega_{c}(\gamma_{1}+\gamma_{2}+1)$.
It follows that, when the initial state is fully contained in the LLL, both the center-of-mass wave function and relative distance wave function will be eigenstates in the LLL. For density-density interactions only the relative distance wave function evolves, while the center-of-mass wave function is still an eigenstate of the Hamiltonian. This implies that, even in the presence of interband transitions, as long as the initial center of mass component of the wavefunction lives in a single LL, the real space dipole in the $x$-direction is conserved by the dynamics. Thus, both average center of mass and its higher-moments are constant, and the dipole moment in real space is preserved to all orders. This reasoning can be directly generalized to more particles as well.
@@
When the center-of-mass wave function of the initial state is not an eigenstate, i.e., when the initial state of the center of mass is not confined in the LLL, higher moments of the dipole moment will not be conserved in general. We consider an initial state that is a superposition of different Landau levels both in the center of mass and relative coordinate $\Psi(\mathbb{R},\mathbb{r})=\sum_{\gamma\beta,n,m}c_{\gamma,\beta}^{n,m}\psi_{\gamma}^{n}(\mathbb{R})\phi_{\beta}^{m}(\mathbb{r})$, where $\psi(\mathbb{R})$ and $\phi(\mathbb{r})$ are eigenstates of the non-interacting problem for the center of mass and relative coordinate, respectively. Specifically they are given by
$\psi_{\gamma}^{n}(\mathbb{R})=\frac{1}{\sqrt{{\pi}\gamma!2^{\gamma}L_{y}}}H_{\gamma}\bigg{(}\frac{X-X_{n}}{L_{B}}\bigg{)}e^{-\frac{1}{2}(\frac{X-X_{n}}{L_{b}})^{2}}e^{iK_{n}Y}$ (12)
@@
where $K_{n}=2\pi n/L_{y}$, $X_{n}=L_{b}^{2}K_{n}$, and the magnetic length for the center of mass coordinate $L_{b}=\frac{l_{b}}{\sqrt{2}}$, as well as,
$\phi^{m}_{\beta}(\mathbb{r})=\frac{1}{\sqrt{{\pi}\beta!2^{\beta}L_{y}}}H_{\beta}\bigg{(}\frac{x-x_{m}}{\ell_{b}}\bigg{)}e^{-\frac{1}{2}(\frac{x-x_{m}}{\ell_{b}})^{2}}e^{ik_{m}y}$ (13)
@@
where $k_{m}=2\pi m/L_{y}$, $x_{\beta}=\ell_{b}^{2}k_{m}$, and the magnetic length for the relative coordinate $\ell_{b}=l_{b}\sqrt{2}$. By using the properties of the Hermite polynomials $H_{n}$ we determine the evolution of the moments of the center-of-mass in real-space in the $x$-direction (i.e., dipole moment in the $x$-direction in real space) $\mathcal{P}=X$ and find that for general initial states the center of mass is not conserved.
@@
As a concrete example, we consider an initial state where the two particles have momenta $k_{y}=0$ and that is invariant under inversion symmetry in the spatial coordinate. Since the two particles are bosons the wave function $\Psi(\mathbb{R},\mathbb{r})=\sum_{\alpha,\beta,n,m}c_{\gamma,\beta}^{n,m}\psi_{\gamma}^{n}(\mathbb{R})\phi_{\beta}^{m}(\mathbb{r})$, where $\psi(\mathbb{R})$ and $\phi(\mathbb{r})$ must be invariant under exchange of their coordinates $\Psi(\mathbb{R},-\mathbb{r})=\Psi(\mathbb{R},\mathbb{r})$, which implies $m$ must be even. Considering the state to be invariant under inversion symmetry, $\Psi(-\mathbb{R},-\mathbb{r})=\Psi(\mathbb{R},\mathbb{r})$, $n$ is even as well. Under these assumptions, the average position of the center of mass does not evolve. For concreteness, we show the evolution of the initial state $\Psi_{0}(\mathbb{R},\mathbb{r})=c^{0,0}_{0,0}\psi^{0}_{0}(\mathbb{R})\phi^{0}_{0}(\mathbb{r})+c^{0,0}_{2,2}\psi^{0}_{2}(\mathbb{R})\phi^{0}_{2}(\mathbb{r})$, for different ratios of $|c^{0,0}_{2,2}/c^{0,0}_{0,0}|$ in Fig 4. The fluctuations of the center of mass oscillate with frequency $2\omega_{c}$ when multiple Landau levels are occupied. When the initial state is fully contained in the LLL, $c^{0,0}_{2,2}=0$, the fluctuations of the dipole-moment $\langle\Delta\mathcal{P}^{2}\rangle$ (as well as all higher moments) are constant, as argued before. When creating such an initial state in the LLL with a finite particle density, the subsequent relaxation dynamics in real space will be governed by fracton hydrodynamics; see discussion in the previous section.
@@ IV.2 Synthetic gauge fields for ultracold atoms in optical lattices
Charge neutral atoms in optical lattices with artificial gauge fields are effectively described by the interacting Harper-Hofstadter model, Eq. (2), and geometric conditions on the lattice can be defined such as the non-interacting bands are flat.
@@
We now investigate the full dynamics of this model and initialize the system at time $t=0$ with particles occupying orbitals of the lowest band of the non-interacting problem $H_{0}$. Experimentally, such a state can be obtained by preparing plane-wave states in the $y$-direction, corresponding to an eigenstate of the non-interacting Hofstadter model with ${t_{x}}/{t_{y}}\to 0$, followed by a slow increase of ${t_{x}}/{t_{y}}$; c.f. Ref. [26] where a similar approach has been considered. At time $t=0$ the on-site interaction energy is turned on and the particles start to evolve.
@@
Contrary to the continuum case, the Hamiltonian is not separable in center of mass and relative coordinate degrees of freedom. Thus, although the states exhibit an average conservation of the center of mass along the $x$-direction, $\mathcal{P}=\sum_{x,y}x\,n_{x,y}$, when their density profile is invariant under inversion symmetry (not shown), the fluctuations $\langle\Delta^{2}\mathcal{P}\rangle=\langle\mathcal{P}^{2}\rangle-\langle\mathcal{P}\rangle^{2}$ will not be constant in the presence of interband transitions (even when initially prepared in the LLL). On the other hand, when the dynamics is projected in the LLL, the value of the fluctuations (as well as higher momenta of the center of mass) is preserved. This is because similarly to the continuum the shape of the wave functions in the LLL remain invariant for different momenta and only their wave function center shifts along the $x$-direction, see App. D. Hence, a squeezing motion of the column density is expected in real space provided the dynamics is confined to the LLL. This formally follows also from the fact that the center of mass along the $x$-direction $\mathcal{P}$ projected onto the LLL, commutes with the projected interactions, see App. D. By contrast, for interlevel processes $\mathcal{P}$ is not conserved in real space, despite the total momentum $\sum k_{y}n_{k_{y}}$ being conserved, where the occupation of momentum $k_{y}$, $n_{k_{y}}=\sum_{\gamma}n_{k_{y},\gamma}$, is obtained by summing over all Landau levels $\gamma$. Eigenfunctions with the same $k_{y}$ but on different Landau levels thus have different expectation values of the dipole-moment $\mathcal{P}$ evaluated in real space. We compute numerically the fluctuations of the dipole moment $\langle\Delta^{2}\mathcal{P}\rangle=\langle\mathcal{P}^{2}\rangle-\langle\mathcal{P}\rangle^{2}$ and find that they are reduced with decreasing interactions as the dynamics becomes increasingly confined to the LLL, see Fig. 5(d), consistent with the discussion above. In summary, on the lattice the real-space dipole moment is only conserved when the dynamics is limited to one LL, since interband processes do not preserve dipole moment (center of mass) in real space. Thus, for observables integrated over the y-direction but local in the $x$-direction, a subdiffusive relaxation dynamics will arise only in the limit of small interactions compared to the difference of energies between different bands.
@@ V Discussion and Outlook
Recent experiments with ultracold atoms have realized models with synthetic gauge fields. In these systems, non-equilibrium quantum dynamics are most naturally accessible. For this reason, it is crucial to understand relaxation processes for states with partially filled Landau levels.
In contrast with the common anticipation that interacting quantum many-body systems relax diffusively, we showed that this is not the case for finite energy density states in the lowest Landau level. Instead, the relaxation dynamics of finite energy density states confined within a Landau level effectively occurs in a one-dimensional lattice and is constrained by dipole conservation in the effective lattice. Hence, the far-from-equilbrium dynamics in the LLL are governed by fracton hydrodynamics, characterized by a slow subdiffusive relaxation of density excitations and their fluctuations. The properties of the relaxation are tunable by the effective thickness of the torus. The equivalence of dipole conservation in the LLL and dipole conservation in real space along the $x$-direction allows for a direct measurement of the unconventional relaxation dynamics, both in rotating quantum gases and in interacting Harper-Hofstadter models with certain geometric constraints.
For future work, it will be interesting to analyze the short time dynamics of specific initial states in more detail. Furthermore, one could study the consequences of non-flat bands in the interacting Harper-Hofstadter model, that arise for example with open boundary conditions, on the effective relaxation dynamics. Moreover, excitations of the FQH system could be studied in the low-temperature regime and related to the different dynamical regimes of fracton excitations [66, 67]. It would be also exciting to explore relaxation processes in the vortex dynamics of rotating quantum gases [8]. Previous works have already interpreted vortices as fractons and discussed Hilbert-space fragmentation phenomena and fractonic dynamics in vortex systems [68].
@@ Appendix B Generalization to different geometries and non-flat bands
The analysis in the main text focuses on the flat band condition, obtained directly for rotating quantum gases or for specific choices of lattice geometries in the Harper-Hofstadter model. Here, we generalize the considerations for lattice systems to non-flat bands. Non flat-bands arise in a lattice when the condition $N_{x}=Q=\alpha N_{y}$, is not satisfied and when the boundary conditions are not met. However, for cases where the width of the band $\Delta$ is small compared to the other energy scales, $\Delta\ll U\ll\hbar\omega_{c}$, our expectation is that this will not modify the thermalization properties in the LLL.